A box contains 10 balls. Three are numbered 0, two are numbered 1, one is numbered 2, two are numbered 3, and two are numbered 4. The balls are mixed and one is selected at random. After a ball is selected, its number is recorded. Then it is replaced. If the experiment is repeated many times, find the standard deviation of the numbers on the balls. (Round off your final answer to 2 decimal places).


1
Expert's answer
2022-01-24T17:03:12-0500

Let X be a random variable represents the number on the ball, then

"P(X=0)=0.3"

"P(X=1)=0.2"

"P(X=2)=0.1"

"P(X=3)=0.2"

"P(X=4)=0.2"

"\sigma=\sqrt{D}=\sqrt{M(X^2)-M^2(X)}"

"M(X)=0*0.3+1*0.2+2*0.1+3*0.2+4*0.2=1.8\implies M^2(X)=3.24"

"M(X^2)=0^2*0.3+1^2*0.2+2^2*0.1+3^2*0.2+4^2*0.2=5.6"

So, "\sigma=\sqrt{5.6-3.24}=\sqrt{2.36}\approx1.54"